# EditCtrl: Disentangled Local and Global Control for Real-Time Generative Video Editing

### v1

@@ 1 Introduction
In this paper, we address this critical efficiency bottleneck and propose a novel framework, EditCtrl, that decouples localized sparse inpainting generation from the global video context, enabling high-fidelity generation results at a fraction of the computational cost. Instead of processing the entire video, our approach strategically allocates compute only to the pixels that require editing. Our approach achieves this with two key components: i) a sparse local context module that operates exclusively on tokens within the target edit masks; and ii) a lightweight temporal global context embedder that captures video-wide coherence and guides the local generation process.
This design yields several transformative advantages. First, the computational cost becomes proportional to the size of the target edit mask, independent of the input video resolution. Second, our adapter modules are added without updating the model weights of the base diffusion model, preserving its powerful generative priors. This design, unlike previous methods that fine-tune the base video diffusion model, makes EditCtrl natively compatible with distilled autoregressive models for content propagation, further accelerating inference latencies.
@@ 2 Related Works
Early data-driven video inpainting methods used spatiotemporal transformers or convolutional networks to attend to the global context of the video while propagating information across frames using optical flow (zhou2023propainter; E2FGVI; zhang2024blur; Zhang_2022_FGT). These methods efficiently inpaint target regions, but are limited in their ability to handle complex motion, masked regions, or generate coherent semantic content given a condition like a text prompt. To increase video inpainting quality and diversity, generative-based methods repurposed powerful priors in the form of video diffusion models (bian2025videopainteranylengthvideoinpainting; vace; li2025diffueraserdiffusionmodelvideo; liu2024generativevideopropagation; samuel2025omnimattezero; ma2025follow; gen-omnimatte). While generative-based models inpaint high quality content coherently, they take both the entire video as context and a per-pixel mask indicating whether to preserve or generate each pixel. This approach is inefficient for interactive editing since the underlying diffusion model needs to repeatedly generate pixels outside the local masked edit area. Furthermore, it ties mask information to the video context, preventing generative inpainting from making multiple distinct edits at different regions or propagating inpainting information when video frames are not available in tasks like image-to-video generation.
@@
To accelerate video editing, a straightforward approach is to reduce the number of diffusion sampling steps the base video diffusion model needs to produce coherent outputs. This includes knowledge distillation (yin2024onestep; yin2025causvid; luo2023latent; heek2024multistepconsistencymodels), sparse attention (liu2025mardini; zhang2025vsa; shin2025motionstream; zhang2025spargeattn), or linear transformers (xie2024sana; chen2025sana). This can be naively applied to the base video diffusion model used for generative editing, accelerating editing as a result. However, this only reduces the number of underlying diffusion steps and not the data processed for content editing in terms of the local and global context sizes. As such, runtime and memory scale with the input video resolution regardless of the target edit area. Token merging based approaches (bolya2022tome; li2023vidtome; object_centric_video_diffsion) address this by pruning tokens outside the local context, though this leads to significant quality degradation and may actually slow inference due to the token importance calculation step. Some image generation methods compress the global context together with the local context (lazy_diffusion; beyer2025highly), but this requires training specialized generative models that can take the new information, which is expensive and not competitive with the quality of pretrained full-attention diffusion models. This design also prevents integration with additional control variants without modifying the base model weights. Overall, methods that accelerate the base video diffusion model do not target the bottleneck of local-global context for editing and must process the entire video. Alternatively, token-merging and compression approaches are destructive, and runtime may not cleanly scale with the size of the edit. For interactive editing, the computational cost must be decoupled from the global context of the video and thus the total video resolution.
@@
EditCtrl separates the low and high frequency components of the mask, which correspond to the global and sparse local contexts, to significantly accelerate video editing while matching the quality when using the full global context, making it invariant to video resolution and enabling fast editing. Our approach naturally extends to interactive editing capabilities such as multi-prompt editing for different regions and content propagation, which are challenging for non-disentangled frameworks that rely on full-attention.
@@ 3.1 Preliminaries and Notation
Text-to-video diffusion models are primarily structured as latent diffusion models using a powerful diffusion transformer (DiT) backbone $\epsilon_{\theta}$ to generate visually appealing, temporally consistent content given a text prompt $\mathbf{p}$. The model employs a pretrained variational autoencoder (VAE) consisting of an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ to encode videos to a smaller latent space and reduce computational cost. At training, $\mathcal{E}$ encodes a source video $\mathbf{V}_{\text{src}}\in\mathbb{R}^{F\times H\times W}$, defined as a sequence of $F$ frames of size $H\times W$, into latent space: $\textbf{z}=\mathcal{E}(\mathbf{V}_{\text{src}})$. A sampled noise $\epsilon_{t}$ is added to the encoded latent using a scheduler (ho2020ddpm) with timestep $t\in\{1,\cdots,T\}$. $\epsilon_{\theta}$ takes the produced noisy video latent $\mathbf{z}^{t}$ to predict the noise $\epsilon_{t}$ in the forward diffusion process:
$\mathcal{L}_{\text{DM}}=\|\epsilon_{\theta}\left(\mathbf{z}^{t},t;\mathbf{p}\right)-\epsilon_{t}\|_{2}^{2},$ (1)
To make use of the generative prior knowledge in pretrained text-to-video diffusion models for video editing, we train a conditional model $c_{\phi}$ to encode additional conditional information $\mathbf{C}$ and steer the predictions from $\epsilon_{\theta}$. $\mathbf{C}$ is derived from editing mask frames $\mathbf{V}_{m}\in\mathbb{R}^{F\times H\times W}$ corresponding to target edit regions. These are used to mask the foreground in the input video, giving the background video $\mathbf{V}_{b}\in\mathbb{R}^{F\times H\times W}$ that is encoded with $\mathcal{E}$. $\mathbf{C}$ is constructed as a channel-wise concatenation of the information $(\mathcal{E}(\mathbf{V}_{b}),\mathbf{V}_{m}^{\downarrow})$, where $\mathbf{V}_{m}^{\downarrow}$} are the editing masks that are down-sampled to the latent resolution, which gives our conditional diffusion process:
$\mathcal{L}_{\text{CDM}}=\|\epsilon_{\theta}\left(\mathbf{z}^{t},t;\mathbf{p},c_{\phi}(\mathbf{C})\right)-\epsilon_{t}\|_{2}^{2},$ (2)
@@ 3.2 Disentangled Editing Architecture Design
Our goal is to design an architecture that generates video content in a desired region while efficiently incorporating global temporal context from the rest of the input video. EditCtrl builds on a ControlNet-like architecture (zhang2023controlnet), with a trainable context control module used to guide a frozen, pre-trained video diffusion model (HaCohen2024LTXVideo; vace; nvidia2025cosmostransfer1conditionalworldgeneration). This context module is trained for full-frame inpainting with bi-directional full-attention, processing the entire video $\mathbf{V}_{\text{src}}$ and editing masks $\mathbf{V}_{m}$, making it computationally expensive regardless of the edit size.
Our key architectural insight is to disentangle this process. We modify the framework by introducing two control adapters: i) a local context adapter module re-purposed from the original control by fine-tuning it to operate only on tokens within the masked region (Sec. 3.2.1); and ii) a new, lightweight temporal global context embedder that efficiently encodes the full video context to maintain global coherence (Sec. 3.2.2). This design preserves the base model’s high-quality generative capabilities while making the computational cost of editing proportional to the target area size, not the full video resolution.
3.2.1 Local Context Encoder
To repurpose the inpainting control module for local context editing, we use $\mathbf{V}_{m}^{\downarrow}$ as an attention mask for the tokens corresponding to the foreground. Background tokens $\mathbf{C}$ are selected with $\mathbf{V}_{m}^{\downarrow}$ to produce $\mathbf{C}_{\text{local}}$. Masking out all the background tokens results in poor blending, so $\mathbf{V}_{m}^{\downarrow}$ is dilated to neighboring pixels before selection (Liu_2021_FuseFormer). This attention mask operation is also applied to $\mathbf{z}^{t}$, and the output from the control module is summed element-wise with the feed-forward network output after selected video transformer layers. By processing only the corresponding local tokens through the video diffusion model and the control module, the diffusion process is proportionally accelerated. However, generation quality is much worse as the pretrained control module was trained with full-attention and not sparse attention patterns corresponding to the target mask region. Thus, the module is fine-tuned into a local context encoder using a mask-aware diffusion loss
@@ 3.2.1 Local Context Encoder
3.2.2 Global Context Embedder
The local context encoder could generate high-fidelity content from local context tokens alone, but it suffers from having no access to the surrounding context. Without the global temporal context of the video, the local encoder cannot integrate important structure into the generation, such as appearance and scene cues (e.g., lighting, structure, dynamics, camera motion). To ensure that local edits are consistent with the full video context, we insert a lightweight global context embedder module $G_{\psi}$ into the base video model to integrate global information during diffusion.
@@ 3.2.2 Global Context Embedder
To incorporate global information at generation, we enrich the cross-attended features $\mathbf{x}$ produced with the video DiT by injecting temporally-aware attention features as shown in Fig. 3. This is similar to using CLIP for image-guided generation, but in our case we use a temporal global context representation computed using the VAE encoder $\mathcal{E}(\mathbf{V}_{b}^{\downarrow})$, denoted as $\mathbf{z}_{b}$. $\mathbf{V}_{b}$ is spatially down-sampled to a resolution of $256\times 256$ regardless of the original resolution to give $\mathbf{V}_{b}^{\downarrow}$. This is to increase robustness to aspect ratio and the number of frames we can encode. $\mathbf{z}_{b}$ is then processed by a trainable patch layer to produce the global context token embeddings, efficiently capturing video-wide temporal evolution and high level scene cues while remaining invariant to the original video resolution. $G_{\psi}$ takes the global context embeddings to calculate the attention weights of query tokens $Q$ with the global feature keys and values $\mathbf{K}_{g},\mathbf{V}_{g}$. These are added to the features after cross-attending to the text prompt
@@
where $\mathbf{W}_{0}$ is a zero initialized linear layer. This attention operation enables minimal yet sufficient control over the text prompt embedding, enriching the features without degrading the information from the text embedding. $G_{\psi}$ thus steers the generation of local token features to ensure they align with the text prompt and high level global video context with minimal compute overhead. We define the mask-aware loss function with $\mathbf{z}_{b}$ as
$\mathcal{L}_{\psi}=\|\epsilon_{\theta}\left(\mathbf{z}^{t},t;\mathbf{p},G_{\psi}(\mathbf{z}_{b}),c_{\phi}(\mathbf{C}_{\text{local}})\right)-\epsilon_{t}\odot\mathbf{V}_{m}^{\downarrow}\|_{2}^{2},$ (5)
@@ Appendix A Additional Visualizations
We show additional results editing in distinct regions with multiple prompts as well as content propagation in Figs. 8-9. Furthermore, we provide additional visualizations of EditCtrl’s video editing and inpainting results on VPBench and DAVIS in Figs. 10-13.

### v2

@@ 1 Introduction
We are inspired by prior work that overcomes this bottleneck by efficiently allocating computation for inpainting using tokens inside the target edit mask region (zhou2023propainter; lazy_diffusion). In particular, this principle can be applied in generative image editing by finetuning a pre-trained diffusion model to only focus on relevant local features while using a compressed global representation as context (lazy_diffusion). We extend this approach to tackle the challenges in generative video editing: understanding and inpainting temporally coherent content, and enabling intelligent video-specific downstream operations like content propagation. Specifically, we seek to leverage high-quality video generation models in a disentangled manner without harming the pre-trained quality and achieve this via learning lightweight adapter modules.
In this paper, we propose EditCtrl, a disentangled generative video editing framework for temporally coherent video inpainting. Our design follows two principles: i) computation should scale with the edit region, not the video resolution; and ii) conditioning should be injected via adapters on a frozen base model, preserving its generative quality and compatibility with model variants. We instantiate these through a sparse local context encoder that processes only tokens within the edit mask, and a lightweight temporal global context embedder that provides video-wide cues to guide local generation.
Because EditCtrl achieves proportional compute scaling entirely through lightweight adapters on a frozen base model, it preserves the full generative quality of the pretrained backbone while grounding local generation in global temporal context for coherent edits. This non-destructive design makes EditCtrl natively compatible with model variants such as distilled and autoregressive models, unlocking downstream intelligent video editing applications, including multi-region editing with distinct prompts and real-time content propagation into future frames.
@@ 2 Related Works
Early data-driven video inpainting methods used spatiotemporal transformers or convolutional networks to attend to the global context of the video while propagating information across frames using optical flow (zhou2023propainter; E2FGVI; zhang2024blur; Zhang_2022_FGT). These methods efficiently inpaint target regions, but are limited in their ability to handle complex motion, masked regions, or generate coherent semantic content given a condition like a text prompt. To increase video inpainting quality and diversity, generative-based methods repurposed powerful priors in the form of video diffusion models (bian2025videopainteranylengthvideoinpainting; vace; li2025diffueraserdiffusionmodelvideo; liu2024generativevideopropagation; samuel2025omnimattezero; ma2025follow; gen-omnimatte). While generative-based models inpaint high-quality content coherently, they take both the entire video as context and a per-pixel mask indicating whether to preserve or generate each pixel. This approach is inefficient for interactive editing since the underlying diffusion model needs to repeatedly generate pixels outside the local masked edit area. Furthermore, it ties mask information to the video context, preventing generative inpainting from making multiple distinct edits at different regions or propagating inpainting information when video frames are not available in tasks like image-to-video generation.
@@
To accelerate video editing, a straightforward approach is to reduce the number of diffusion sampling steps the base video diffusion model needs to produce coherent outputs. This includes knowledge distillation (yin2024onestep; yin2025causvid; luo2023latent; heek2024multistepconsistencymodels), sparse attention (liu2025mardini; zhang2025vsa; shin2025motionstream; zhang2025spargeattn), or linear transformers (xie2024sana; chen2025sana). This can be naively applied to the base video diffusion model used for generative editing, accelerating editing as a result. However, this only reduces the number of underlying diffusion steps and not the data processed for content editing in terms of the local and global context sizes. As such, runtime and memory scale with the input video resolution regardless of the target edit area. Token merging based approaches (bolya2022tome; li2023vidtome; object_centric_video_diffsion) address this by pruning tokens outside the local context, though this leads to significant quality degradation and may actually slow inference due to the token importance calculation step. Some generation methods compress the global context together with the local context (beyer2025highly), but this requires training specialized generative models that can take the new information, which is expensive and not competitive with the quality of pretrained full-attention diffusion models. We instead adopt insights from an efficient image editing approach (lazy_diffusion) that fine-tunes a pretrained diffusion model to only operate on local tokens with additional compressed global context, thus accelerating diffusion proportionally relative to the masked region area.
Our work extends this insight to the video domain while only learning local adapters instead of full finetuning, thus allowing integration with additional control variants without modifying the base model weights.
@@
EditCtrl uses lightweight adapters to disentangle the low and high frequency components of the mask, which correspond to the global and sparse local contexts, to significantly accelerate video editing with minimal interference in the base model. By keeping the base model frozen, our approach remains invariant to video resolution, enables fast editing, and remains compatible with additional control variants. It naturally extends to interactive video editing capabilities such as multi-prompt editing for different regions and temporal content propagation, which are challenging for non-disentangled frameworks that rely on full-attention.
@@ 3.1 Preliminaries and Notation
Text-to-Video Diffusion.
Text-to-video diffusion models are primarily structured as latent diffusion models using a powerful diffusion transformer (DiT) backbone $\epsilon_{\theta}$ to generate visually appealing, temporally consistent content given a text prompt $\mathbf{p}$. The model employs a pretrained variational autoencoder (VAE) consisting of an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ to encode videos to a smaller latent space and reduce computational cost. At training, $\mathcal{E}$ encodes a source video $\mathbf{V}_{\text{src}}\in\mathbb{R}^{F\times H\times W}$, defined as a sequence of $F$ frames of size $H\times W$, into latent space: $\mathbf{z}=\mathcal{E}(\mathbf{V}_{\text{src}})$. A sampled noise $\epsilon_{t}$ is added to the encoded latent using a scheduler (ho2020ddpm) with timestep $t\in\{1,\cdots,T\}$. $\epsilon_{\theta}$ takes the produced noisy video latent $\mathbf{\bar{z}}^{t}$ to predict the noise $\epsilon_{t}$ in the forward diffusion process:
$\mathcal{L}_{\text{DM}}=\|\epsilon_{\theta}\left(\mathbf{\bar{z}}^{t},t;\mathbf{p}\right)-\epsilon_{t}\|_{2}^{2},$ (1)
Control Adapters.
To make use of the generative prior knowledge in pretrained text-to-video diffusion models for video editing, a conditional adapter module $c_{\phi}$ is trained to compress additional conditional information $\mathbf{C}$ and steer the predictions from $\epsilon_{\theta}$. $\mathbf{C}$ is derived from editing mask frames $\mathbf{V}_{m}\in\mathbb{R}^{F\times H\times W}$ corresponding to target edit regions. These are used to mask the foreground in the input video, giving the background video $\mathbf{V}_{b}\in\mathbb{R}^{F\times H\times W}$ that is encoded with $\mathcal{E}$. $\mathbf{C}$ is constructed as a channel-wise concatenation of the information $(\mathcal{E}(\mathbf{V}_{b}),\mathbf{V}_{m}^{\downarrow})$, where $\mathbf{V}_{m}^{\downarrow}$} are the editing masks that are down-sampled to the latent resolution, which gives our conditional diffusion process:
$\mathcal{L}_{\text{CDM}}=\|\epsilon_{\theta}\left(\mathbf{\bar{z}}^{t},t;\mathbf{p},c_{\phi}(\mathbf{C})\right)-\epsilon_{t}\|_{2}^{2},$ (2)
Context-Guided Local Computation.
LazyDiffusion (lazy_diffusion) introduced a principle for diffusion-based efficient (image) editing where computation is restricted to the edit region and edit context can be supplied through a compact side channel. A diffusion model is fine-tuned to process noisy latent tokens $\mathbf{z}^{t}$ strictly within the masked region while a separate encoder captures local and global information context $\mathbf{C}_{\text{enc}}$. $\mathbf{C}_{\text{enc}}$ is concatenated to the hidden dimension at every denoising iteration, achieving acceleration proportional to the mask size with global context. A pre-trained diffusion model is finetuned with a “lazy” diffusion process:
$\mathcal{L}_{\text{Lazy-DM}}=\|\epsilon_{\theta}\left(\mathbf{z}^{t}\oplus\mathbf{C}_{\text{enc}},t;\mathbf{p}\right)-\epsilon_{t}\odot\mathbf{V}_{m}^{\downarrow}\|_{2}^{2},$ (3)
@@ 3.2 Disentangled Editing Architecture Design
Our goal is to design an efficient video editing framework that scales proportionally with respect to edit area and flexibly composes with adapters that were pre-trained for full-attention tasks. Towards this, we build on the design choices highlighted in Sec. 3.1 of achieving flexibility with control adapters and efficiency via context-guided local computation. Specifically, we i) disentangle video editing with local encoder and global embedder adapters that separately operate on local context tokens in addition to efficiently encoding the full global temporal context to maintain coherence, and ii) achieve this with adapter-based finetuning. This design combination focuses compute in masked regions with temporal coherence, preserves the base model’s generative capabilities, and enables intelligent downstream video editing applications.
Context-Guided Video Diffusion
EditCtrl builds on a ControlNet-like architecture (zhang2023controlnet), with a trainable context control module used to guide a frozen, pre-trained video diffusion model (HaCohen2024LTXVideo; vace; nvidia2025cosmostransfer1conditionalworldgeneration). This context module is trained for full-frame inpainting with bi-directional full-attention, processing the entire video $\mathbf{V}_{\text{src}}$ and editing masks $\mathbf{V}_{m}$, making it computationally expensive regardless of the edit size. To repurpose the inpainting control module for local context editing, we use $\mathbf{V}_{m}^{\downarrow}$ as an attention mask for the tokens corresponding to the foreground. Background tokens $\mathbf{C}$ are selected with $\mathbf{V}_{m}^{\downarrow}$ to produce $\mathbf{C}_{\text{local}}$. Masking out all the background tokens results in poor blending, so $\mathbf{V}_{m}^{\downarrow}$ is dilated to neighboring pixels before selection (Liu_2021_FuseFormer). This attention mask operation is also applied to $\mathbf{z}^{t}$, and the output from the control module is summed element-wise with the feed-forward network output after selected video transformer layers. By processing only the corresponding local tokens through the video diffusion model and the control module, the diffusion process is proportionally accelerated. However, generation quality is much worse as the pretrained control module was trained with full-attention and not sparse attention patterns corresponding to the target mask region. Thus, the module is fine-tuned into a local context encoder using a mask-aware diffusion loss
@@
To incorporate global information such as appearance and scene cues (e.g., lighting, structure, dynamics, camera motion) at generation, we enrich the cross-attended features $\mathbf{x}$ produced with the video DiT by injecting temporally-aware attention features as shown in Fig. 3. This is similar to using CLIP for image-guided generation, but in our case we use a temporal global context representation computed using the VAE encoder $\mathcal{E}(\mathbf{V}_{b}^{\downarrow})$, denoted as $\mathbf{C}_{\text{global}}$. $\mathbf{V}_{b}$ is spatially down-sampled to a resolution of $256\times 256$ regardless of the original resolution to give $\mathbf{V}_{b}^{\downarrow}$. This is to increase robustness to aspect ratio and the number of frames we can encode. $\mathbf{C}_{\text{global}}$ is then processed by a trainable patch layer to produce the global context token embeddings, efficiently capturing video-wide temporal evolution and high level scene cues while remaining invariant to the original video resolution. The global embedder $G_{\psi}$ takes the global context embeddings to calculate the attention weights of query tokens $Q$ with the global feature keys and values $\mathbf{K}_{g},\mathbf{V}_{g}$. These are added to the features after cross-attending to the text prompt
@@
where $\mathbf{W}_{0}$ is a zero initialized linear layer. This attention operation enables minimal yet sufficient control over the text prompt embedding, enriching the features without degrading the information from the text embedding. $G_{\psi}$ thus steers the generation of local token features to ensure they align with the text prompt and high level global video context with minimal compute overhead. We define the mask-aware loss function with $\mathbf{C}_{\text{global}}$ as
$\mathcal{L}_{\psi}=\|\epsilon_{\theta}\left(\mathbf{z}^{t},t;\mathbf{p},G_{\psi}(\mathbf{C}_{\text{global}}),c_{\phi}(\mathbf{C}_{\text{local}})\right)-\epsilon_{t}\odot\mathbf{V}_{m}^{\downarrow}\|_{2}^{2},$ (6)
@@ Appendix A Additional Visualizations
We show additional results editing in distinct regions with multiple prompts as well as content propagation in Figs. 10-11. Furthermore, we provide additional visualizations of EditCtrl’s video editing and inpainting results on VPBench and DAVIS in Figs. 12-15.
@@ Appendix B Additional Details
$G_{\psi}$ Resolution & Transferability.
To test the fixed $\times256$ global embedder resolution and transferability, we test EditCtrl on a challenging 21:9 ultrawide timelapse video by itself and with a publicly available aesthetic style LoRA (diffsynth_studio_wan2.1_1.3b_lora_aesthetics_v1) in Fig. 7. This example demonstrates that the downsampled global context captures sufficient spatial and temporal information even for extreme aspect ratios and complex scenes, and moreover, that our adapter design composes seamlessly with style LoRAs, confirming the non-destructive design generalizes beyond the base model. This is shown by the enhanced style EditCtrl generates when using an aesthetic LoRA.
Failure Modes.
We present two cases of inputs where EditCtrl fails in Fig. 8. In the first case on the left, at high motion EditCtrl introduces artifacts and incorrect interaction with the surrounding scene. The artifacts are also present for the full-attention baseline VACE, indicating that this is a structural problem that we attribute to the VAE. In the second case on the right, as the mask becomes exceedingly large finegrained editing becomes more difficult and EditCtrl changes the content by more than desirable.
Figure 7: Editing complex timelapse 21:9 ultrawide footage. EditCtrl can also be used with additional conditioning LoRAs (diffsynth_studio_wan2.1_1.3b_lora_aesthetics_v1).
Figure 8: Failure Modes for EditCtrl on High Motion (a) & Inaccurate Masks (b).